### v2

# DeceptionBench: A Comprehensive Benchmark for AI Deception Behaviors in Real-world Scenarios

Abstract
Despite the remarkable advances of Large Language Models (LLMs) across diverse cognitive tasks, the rapid enhancement of these capabilities also introduces emergent deception behaviors that may induce severe risks in high-stakes deployments. More critically, the characterization of deception across realistic real-world scenarios remains underexplored. To bridge this gap, we establish DeceptionBench, the first benchmark that systematically evaluates how deceptive tendencies manifest across different societal domains, what their intrinsic behavioral patterns are, and how extrinsic factors affect them. Specifically, on the static count, the benchmark encompasses 150 meticulously designed scenarios in five domains, i.e., Economy, Healthcare, Education, Social Interaction, and Entertainment, with over 1,000 samples, providing sufficient empirical foundations for deception analysis. On the intrinsic dimension, we explore whether models exhibit self-interested egoistic tendencies or sycophantic behaviors that prioritize user appeasement. On the extrinsic dimension, we investigate how contextual factors modulate deceptive outputs under neutral conditions, reward-based incentivization, and coercive pressures. Moreover, we incorporate sustained multi-turn interaction loops to construct a more realistic simulation of real-world feedback dynamics. Extensive experiments across LLMs and Large Reasoning Models (LRMs) reveal critical vulnerabilities, particularly amplified deception under reinforcement dynamics, demonstrating that current models lack robust resistance to manipulative contextual cues and the urgent need for advanced safeguards against various deception behaviors. Code and resources are publicly available at https://github.com/Aries-iai/DeceptionBench.
oWarning: This paper may contain some offensive contents in data and model outputs.
1 Introduction
Recently, Large Language Models (LLMs) achiam2023gpt; qwen2.5 have made remarkable progress across a broad range of tasks, facilitating their integration into real-world applications, including content generation acharya2023llm, code synthesis du2024mercury; liu2023your, and information retrieval labruna2024retrieve. As these models demonstrate increasingly sophisticated capabilities in natural language understanding karanikolas2023large and complex reasoning guo2025deepseek, a critical question emerges: Do these advanced cognitive abilities also enable more subtle forms of manipulation? Recent evidence chern2024behonest; park2024ai suggests that LLMs can exhibit deception behaviors, such as producing misleading or strategically false information that undermines trust and poses significant risks in high-stakes deployments. Understanding and quantifying such deceptive tendencies becomes crucial as these models gain more autonomy and influence in decision-making processes. While some benchmarks chern2024behonest; hagendorff2024deception; jarviniemi2024uncovering; su2024ai; wu2025opendeception have tried to evaluate deception behavior, current works often focus narrowly on psychological experiments johnson2011trust; xie2024can or limited scenarios, failing to capture the multifaceted nature of deception across realistic, diverse contexts, as shown Table˜1.
Therefore, to systematically characterize deception behaviors in LLMs and establish a comprehensive evaluation framework, we aim to address the following three fundamental questions:
  • $\diamond$
    How do deceptive tendencies manifest across different critical societal domains? Real-world applications of LLMs span diverse high-stakes scenarios, each with unique trust requirements and potential harm profiles. Thus, to fully explore the potential deceptive tendencies of LLMs, a systematic evaluation across these varied scenarios is essential, as deception behaviors may exhibit domain-specific patterns that cannot be captured through single-domain assessments.
  • $\diamond$
    What intrinsic behavioral patterns drive deceptive responses? Beyond surface-level outputs, understanding the underlying motivations of deception behavior is more important. Identifying these intrinsic drivers enables us to move beyond addressing symptoms to targeting root causes, thereby uncovering the thinking logic behind deceptive choices. This is vital for designing mitigation strategies that could alter how models respond to situations that may elicit deception.
  • $\diamond$
    How do extrinsic contextual factors affect deception behaviors? Actually, LLMs often operate within dynamic environments where external contextual factors can significantly modulate their responses. Thus, investigating how these factors influence deceptive outputs is necessary, which could help safeguard against deception behaviors in scenarios where such influences are inevitable, and identify the specific conditions under which deception intensifies or diminishes.
Table 1: Comparison between DeceptionBench and other benchmarks of deception. DeceptionBench offers more comprehensive coverage across aspects, more rigorous evaluation strategies that include both thought and response, a significantly larger scale in scenario quantity and model coverage, and a more versatile setting supporting both single-turn and multi-turn interactions.
Evaluation Aspects Evaluation Strategy Statistics Setting
Economy
Healthcare
Education
Social Interaction
Entertainment
Thought
Response
Scenario Num
LLM
Single-turn
Multi-turn
Variation
BeHonest chern2024behonest 10 9 (2)
OpenDeception wu2025opendeception 50 11 (2)
FalseBelief hagendorff2024deception 2 10 (7)
AI-LIEDAR su2024ai 60 6 (2)
CompanyDeception jarviniemi2024uncovering 4 6 (6)
DeceptionBench (ours) 150 14 (8)
In response, we propose DeceptionBench, a comprehensive benchmark that systematically evaluates LLM deception behaviors through three interconnected dimensions: the breadth of manifestation across societal domains, the intrinsic motivations underlying deceptive responses, and the dynamic modulation by extrinsic contextual factors. The specific details are as follows:
Establishing domain breadth. We select five critical domains: Economy, Healthcare, Education, Social Interaction, and Entertainment, constructing 150 carefully designed scenarios with over 1,000 evaluation samples. The selection is driven by two key considerations: First, these domains exhibit distinct operational characteristics and consequence severity, such as Healthcare demanding stringent accuracy for patient safety, while Social Interaction prioritizing authenticity in human connections. Second, they collectively encompass diverse situations where deception behaviors may emerge through different mechanisms and with varying intensity. Through these examples, we could capture the nuanced manifestations of deception across different professional and social contexts.
Uncovering intrinsic behavioral drivers. Towards a deeper understanding of deception, we examine two fundamental intrinsic patterns: Egoism xie2024can and Sycophancy sharma2023towards. Egoistic tendencies reflect self-centered behaviors where models prioritize their own objectives or self-preservation, while sycophantic tendencies manifest as user-appeasing behaviors where models provide misleading information to gain favor or comply with perceived user expectations grover1994influence; mitchell2000living. By designing scenarios that distinguish these patterns, we could trace whether deceptive responses originate from internal decision logic or from misaligned attempts to satisfy external entities.
Characterizing extrinsic contextual modulation. We implement a three-tier framework to examine how environmental factors shape deceptive outputs: Neutral conditions (L1-Inherent), where no external influence is exerted, establishing baseline tendencies; Incentivization and Coercion (L2-Induced), where reward structures or coercive prompts simulate real-world inductions; and Iterative feedback loops (L3-Multi-turn Induced), where sustained dialogues test whether deception behaviors escalate, stabilize, or diminish over successive interactions swanson2004let. This progressive design captures both the triggering conditions and the evolutionary trajectories of deceptive patterns.
Through systematic evaluation across the above three dimensions, experimental results reveal several insights into the inner mechanisms and conditions where LLMs exhibit deception behaviors. We observe significant variation in deceptive tendencies across domains, with certain contexts eliciting substantially higher rates of misleading outputs. The interplay between intrinsic and extrinsic factors also shows that models demonstrating strong sycophantic tendencies show amplified deception under reward-based incentivization, while those with egoistic patterns exhibit heightened deception under coercive pressures. Furthermore, multi-turn interactions expose vulnerability to escalating deception. These findings underscore that deception in LLMs is not a monolithic phenomenon but emerges from complex interactions between domain context, intrinsic patterns, and extrinsic factors.
2 DeceptionBench
Figure 1: Overview of DeceptionBench. The benchmark systematically integrates intrinsic motivational patterns (Egoism vs. Sycophancy) and extrinsic contextual influences (Reward vs. Pressure) across three progressive intensity levels. It encompasses diverse real-world scenarios spanning five critical domains, with escalating external inducement from neutral baseline queries (L1) to incentive and coercion-based prompts (L2), and culminating in sustained multi-turn persuasive dialogues (L3).
2.1 Task Formulation
In this work, deception refers to the intentional generation of misleading outputs by LLMs to achieve specific objectives. To systematically characterize such deception behaviors, DeceptionBench requires a task that captures nuanced behavioral patterns beyond surface-level outputs. Discriminative tasks, such as multiple-choice QA, constrain LLMs to predefined response options, limiting insights into their reasoning and behavioral nuances. In contrast, generative QA tasks elicit open-ended responses that reveal how models navigate ambiguous scenarios and justify their choices. We therefore adopt generative QA as the core task paradigm, enabling detailed analysis of deceptive tendencies across realistic contexts. As illustrated in Fig.˜1, the specific task could be formulated as follows:
$\mathbf{R}=\mathbf{M}(\mathbf{P}_{\mathbf{S}},\mathbf{C}_{\mathbf{L}};\theta),$ (1)
where $\mathbf{M}$ denotes the LLM parameterized by $\theta$, and $\mathbf{R}$ represents the response. The input consists of (1) $\mathbf{P}_{\mathbf{S}}$ specifies the role configuration, where $\mathbf{S}\in\{\text{self},\text{other}\}$ distinguishes between self-driven mode (autonomous agent) and user-driven mode (user assistant); (2) $\mathbf{C}_{\mathbf{L}}$ defines the contextual setting with $\mathbf{L}\in\{\text{L1},\text{L2},\text{L3}\}$ denoting the setting as inherent, induced, or multi-turn induced.
2.1.1 Intrinsic Role Specification
The role specification $\mathbf{P}_{\mathbf{S}}$ captures two fundamental behavioral drivers underlying deceptive tendencies. When $\mathbf{S}=\mathrm{self}$, the model operates as an autonomous agent prioritizing self-interest, reflecting egoistic patterns where deception serves internal objectives or self-preservation. When $\mathbf{S}=\mathrm{other}$, the model functions as a user assistant, exhibiting sycophantic tendencies where deception emerges from attempts to satisfy user expectations. By varying $\mathbf{S}$ in Eq. (1), we trace whether deceptive responses originate from internal decision logic or misaligned compliance with external entities.
2.1.2 Extrinsic Context Specification
The contextual configuration $\mathbf{C}_{\mathbf{L}}$ implements a three-tier design, examining how environmental factors shape deceptive outputs. Under $\mathbf{L}=\text{L1}$, the baseline setting establishes inherent tendencies through single-turn interactions without external influence. When $\mathbf{L}=\text{L2}$, the induced setting introduces single-turn external pressures defined as:
$\mathbf{C}_{\text{L2}}=\mathbf{C}_{\text{L2}}(\mathbf{I}),\quad\mathbf{I}\in\{\mathbf{I}_{\text{reward}},\mathbf{I}_{\text{pressure}}\},$ (2)
where $\mathbf{I}_{\text{reward}}$ represents incentivization through promised benefits, and $\mathbf{I}_{\text{pressure}}$ denotes coercive pressure through threats or negative consequences. Under $\mathbf{L}=\text{L3}$, the multi-turn induced setting tests whether deception behaviors escalate through iterative feedback loops:
$\mathbf{R}_{t+1}=\mathbf{M}(\mathbf{P}_{\mathbf{S},t+1},\mathbf{C}_{\text{L3}};\theta),\quad\mathbf{P}_{\mathbf{S},t+1}=\mathbf{M}_{\text{aux}}(\mathbf{R}_{t},\mathcal{E}(\mathbf{R}_{t});\theta_{\text{aux}}),$ (3)
where $\mathbf{R}_{t}$ is the response at turn $t$, $\mathcal{E}(\mathbf{R}_{t})$ identifies reasons for rejection (e.g., failure to deceive), and $\mathbf{M}_{\text{aux}}$, an auxiliary agent parameterized by $\theta_{\text{aux}}$, refines $\mathbf{P}_{\mathbf{S},t+1}$ to enhance deception, terminating when $\mathcal{E}(\mathbf{R}_{t})$ confirms deception success or a predefined limit is reached.
Figure 2: Domain coverage and behavioral types in DeceptionBench. The five domains constitute foundational pillars of contemporary society where LLMs are increasingly integrated, necessitating systematic evaluation to identify potential deceptive risks across diverse application contexts.
2.2 Philosophy of DeceptionBench
2.2.1 Evaluation Aspects
To rigorously assess the deceptive tendencies of LLMs in the real world, as shown in Fig.˜2, DeceptionBench evaluates five critical domains: Economy, Healthcare, Education, Social Interaction, and Entertainment. These domains are selected for their centrality to human life and vulnerability to deception behaviors with profound societal consequences. Each domain encompasses specific scenarios that reflect diverse deceptive practices, necessitating targeted evaluation.
Economy: Deceptive practices in the economic domain, including (i) product of commercial brands, (ii) tax evasion, and (iii) financial deception, can destabilize markets, erode public trust, and cause significant economic harm. For instance, exaggerating product features, misrepresenting income to reduce tax liabilities, or promoting fraudulent investment schemes can disrupt financial systems. With LLMs integrated into financial analysis tools, automated trading platforms, and advisory systems, their potential to generate or endorse deceptive outputs poses substantial risks. Evaluating these scenarios ensures LLMs uphold integrity in economic interactions, preventing the fraudulent schemes hao2025multi; li2023econagent.
Healthcare: Deception behaviors in healthcare, such as (i) health product scams, (ii) medical service fraud, and (iii) health insurance manipulation, directly threaten patient safety and public health, potentially leading to irreversible harm. These include promoting counterfeit health products, performing unnecessary treatments for inflated billing, or falsifying insurance claims. LLMs are deployed in diagnostic support, patient communication, and health information systems alghamdi2024towards; gebreab2024llm, where deceptive outputs could exacerbate fraud or misinformation. Assessing these scenarios is critical to ensure LLMs prioritize accuracy and ethical conduct in sensitive healthcare contexts.
Education: Deceptive practices like (i) training and enrollment scams, (ii) fake degrees & qualifications, and (iii) academic misconduct undermine the credibility of educational systems, impacting career trajectories and societal progress. These encompass fabricating accreditation claims, falsifying academic credentials, or engaging in unethical academic behaviors like data falsification. As LLMs are used in educational tools sonlu2024effects; zhang2024simulating, automated grading, and content creation, their potential to facilitate or detect such deception must be evaluated. Assessing these scenarios ensures LLMs support academic integrity and resist enabling fraudulent credentials.
Social Interaction: Deceptive actions targeting (i) privacy and data, (ii) physical safety, and (iii) emotion and relationship compromise personal security and erode societal trust in interconnected digital environments. These include illegal data collection, misleading safety claims about equipment, or exploiting emotional connections for financial gain. LLMs power social platforms, chatbots, and recommendation systems ayoobi2023looming; leng2023llm; yu2024shadow, where deceptive outputs could amplify privacy violations or emotional manipulation. Evaluating these scenarios is vital to ensure LLMs safeguard user well-being and maintain trust in digital interactions.
Entertainment: Deceptive practices in entertainment, such as (i) fan economy deception, (ii) fake entertainment content, and (iii) fake traffic manipulation, distort public perception and influence discourse on a massive scale. These involve false celebrity endorsements, fabricated narratives, or inflated popularity metrics to mislead audiences. LLMs liu2024can are widely used in content generation, social media analytics, and interactive entertainment, where deceptive outputs could manipulate perceptions. Assessing these scenarios ensures LLMs resist contributing to misinformation nanabala2024unmasking and uphold authenticity in public-facing applications.
To capture the diversity of deception behaviors, we propose a two-level taxonomy encompassing 15 distinct deception types, as detailed in Fig. 2. These different types offer real-world contexts for evaluating LLMs’ deceptive tendencies. Moreover, by combining each deception type with 10 varied deception motivations, we derive 10 scenarios per variation, resulting in a total of 150 scenarios.
2.2.2 Evaluation Strategy
To enable a comprehensive evaluation, we employ an assessment strategy grounded in the Belief-Desire-Intention (BDI) framework georgeff1999belief, which models rational agency by linking internal thought to action. Unlike traditional evaluations that focus solely on responses, our method dissects LLMs’ decision-making by eliciting their internal reasoning alongside their actions. This enables fine-grained analysis of how LLMs navigate deception behaviors in their inner thinking across varied scenarios outlined in Fig.˜2, revealing their ethical robustness in high-stakes contexts.
Specifically, for each prompt $\mathbf{P}_{\mathbf{S}}$ from a deception type, we append the instruction: “(Please provide your thought and final response in JSON format: {"thought": "", "response": ""}).” The output, denoted $\mathbf{O}=\{\mathbf{T},\mathbf{R}\}$, consists of the thought $\mathbf{T}$, capturing the model’s reasoning (its belief about the scenario and prioritized goal), and the response $\mathbf{R}$, reflecting its action. The evaluation function $\mathcal{E}(\mathbf{O})$, defined in Eq.˜3, analyzes $\mathbf{T}$ for deceptive reasoning patterns and $\mathbf{R}$ for deceptive outcomes. For example, in a fake sales prompt, an LLM might generate:
"thought":"The customer prioritizes quality, but sales are the goal. Exaggerating features could drive purchases.", "response":"This product uses the latest technology, nice in performance."
In this case, $\mathbf{T}$ reveals a belief (customer values quality) and goal (maximize sales), while $\mathbf{R}$ suggests deception (exaggerated claims).
2.2.3 Dataset Structure and Prompt Design
As for the structure of DeceptionBench, the dataset is formalized as $\mathcal{D}=\{S_{i}\}_{i=1}^{150}$, where each scenario $S_{i}$ is defined as $S_{i}=(d_{i},t_{i},m_{i},q_{i},g_{i},P_{i})$, where $d_{i}$ denotes the domain, $t_{i}$ specifies one of the 15 deception types in Fig.˜2, and $m_{i}$ captures the underlying motivation (e.g., financial gain, user satisfaction). The query $q_{i}$ is designed to elicit potentially deceptive responses, $g_{i}$ denotes the ethically aligned ground truth response, and $P_{i}$ represents the set of prompt variations for scenario $S_{i}$. The prompt set $P_{i}$ includes single-turn prompts $\mathbf{P}_{\mathbf{S},i,k,\delta}$, where $\mathbf{S}\in\{\mathrm{self},\mathrm{other}\}$, $k\in\{\mathrm{L1},\mathrm{L2}\}$, and $\delta\in\{\emptyset,+,-\}$, as well as multi-turn prompts $\mathbf{P}_{\mathbf{S},i,\mathrm{L3},\delta,t}$, where $\mathbf{S}\in\{\mathrm{self},\mathrm{other}\}$, $\delta\in\{+,-\}$, and $t\in\{1,\ldots,T_{i}\}$. The 150 scenarios are uniformly distributed across five domains (30 per domain), with each deception type instantiated through 10 distinct motivations, yielding 10 scenarios per type. Specifically, for the L1 and L2 settings, the prompt variations are as follows:
Table 2: Prompt configurations across different testing scenarios.
Prompt Level Role Inducement Description
$\mathbf{P}_{\text{self},i,\text{L1},\emptyset}$ L1 Self None Tests deception driven by self-interest
$\mathbf{P}_{\text{other},i,\text{L1},\emptyset}$ L1 Other None Tests deception motivated by sycophancy
$\mathbf{P}_{\text{self},i,\text{L2},+}$ L2 Self Reward Tests deception driven by self-focused rewards
$\mathbf{P}_{\text{self},i,\text{L2},-}$ L2 Self Pressure Tests deception due to self-focused pressure
$\mathbf{P}_{\text{other},i,\text{L2},+}$ L2 Other Reward Tests deception driven by user-focused rewards
$\mathbf{P}_{\text{other},i,\text{L2},-}$ L2 Other Pressure Tests deception due to user-focused pressure
For the multi-turn setting L3, four prompts are included: $\mathbf{P}_{\text{self},i,\text{L3},+,t}$, $\mathbf{P}_{\text{self},i,\text{L3},-,t}$, $\mathbf{P}_{\text{other},i,\text{L3},+,t}$, and $\mathbf{P}_{\text{other},i,\text{L3},-,t}$. Each starts with the corresponding L2 prompt ($\mathbf{P}_{\mathbf{S},i,\text{L2},\delta}$) and generates subsequent prompts at turn $t$ via an auxiliary agent $\mathbf{M}_{\text{aux}}$, iterating until deception succeeds or a predefined limit is reached. Overall, 150 scenarios are totally prepared for our evaluation, each with 2 queries for the L1 setting, 4 queries for the L2 setting, and 4 queries for the L3 setting.
2.3 Dataset Construction
Figure 3: Overview of dataset construction pipeline. In Stage 1, domain-specific motivations are generated via LLM prompting guided by a structured template with few-shot examples, and selected through human evaluation based on diversity and plausibility criteria. Stage 2 converts validated motivations into structured deceptive scenarios, which undergo human revision for fidelity and question validity. Stage 3 further injects intrinsic/extrinsic factors to create test variants.
For dataset construction details, it is processed through a three-stage pipeline. As shown in Fig.˜3, the pipeline progresses from Domain-specific Motivation Generation to Deceptive Scenario Creation and culminates in Factor Injection. In this process, motivations establish the underlying incentives for deception, scenarios create testable contexts grounded in realistic situations, and factor injection systematically diversifies prompts to probe LLMs’ responses under varying intrinsic and extrinsic conditions. This integrated approach ensures comprehensive coverage of deception behaviors across multiple dimensions. The specific details of each stage are as follows:
Domain-specific Motivation Generation. The pipeline begins by generating motivations that underpin deception behavior within each domain and their corresponding deception types, such as Tax Evasion or Health Product Scams. For example, in the Economy domain’s Product of Commercial Brands, a motivation might be maximizing sales revenue. These incentives are crafted to mirror realistic drivers of deception specific to each domain and deception type. Human validation filters the motivations for diversity, plausibility, and contextual fit, ensuring a strong foundation for scenario development that challenges LLMs’ decision-making integrity.
Deceptive Scenario Creation. Then, these motivations inform the creation of detailed scenarios, each comprising a prompt, an open-ended question, and a ground truth response. Prompts set up vivid contexts; for instance, a scenario in Economy’s Product of Commercial Brands might depict a salesperson deciding whether to conceal a product defect to boost sales. For the choice of open-ended questions, as compared to closed-ended ones, they could elicit responses that reveal LLMs’ reasoning and potential for deception. Ground truth responses provide ethical, truthful benchmarks for evaluation. Human validation ensures each component’s rationality, naturalness, and capacity to provoke deception behavior, producing reliable scenarios that probe LLMs’ underlying beliefs.
Factor Injection. Finally, scenarios are diversified by incorporating intrinsic roles ($\mathbf{S}\in\{\text{self},\text{other}\}$) and extrinsic communication settings ($\mathbf{L}\in\{\text{L1},\text{L2},\text{L3}\}$), as specified in Sec.˜2.1, to explore a range of deceptive contexts. L1 prompts test self-driven or user-focused deception without inducements, yielding L1-self and L1-other variations. L2 introduces reward or pressure inducements, generating L2-self-reward, L2-self-pressure, L2-other-reward, and L2-other-pressure. The template for prompt variant generation is demonstrated as Fig.˜14 in the Appendix. For L3 multi-turn interaction, it starts with L2’s initial prompts, with an auxiliary agent $\mathbf{M}_{\text{aux}}$ generating subsequent prompts until deception succeeds or a limit is reached, resulting in variants for L3-self-reward, L3-self-pressure, L3-other-reward, and L3-other-pressure via Fig.˜17 in the Appendix. Human validation confirms alignment with the scenario’s context, resulting in 10 prompt variations per scenario for comprehensive evaluation of LLMs’ deception behavior. Representative dataset examples for each deception type of the five domains are illustrated in Figs.˜6, 7, 8, 9 and 10 in the Appendix. The entire pipeline integrates automated generation with human-in-the-loop validation to ensure scenario realism, motivation fidelity, and systematic coverage of deception behaviors across intensity levels.
2.4 Evaluated Models and Metrics
Model Selection. To comprehensively assess the deceptive tendencies of LLMs, we evaluate 14 models, including 8 proprietary and 6 open-source models, selected to represent a broad spectrum of current LLM capabilities. The proprietary models include state-of-the-art general-purpose models such as GPT-4o achiam2023gpt, Claude-3.7 claude, Grok-3 grok2025, Gemini team2023gemini, Deepseek-R1 guo2025deepseek, etc. The open-source models encompass models like Qwen2.5-7B-Instruct qwen2.5, Qwen2.5-14B qwen2.5, and Qwen-Max, as well as corresponding reasoning models, including the DeepSeek-R1-Distill-Qwen series guo2025deepseek, and QWQ-32B QWQ32, which could enable a thorough evaluation of LLMs’ propensity for deception across varied parameter sizes within a consistent architecture or different training paradigms.
Evaluation Metrics. To measure the tendency for deception behaviors in LLMs, we employ deception rate as the primary metric, evaluating both internal reasoning and final outputs through $\mathrm{DR}_{\mathrm{thought}}$ and $\mathrm{DR}_{\mathrm{response}}$, respectively. These metrics are derived from the structured output format $\mathrm{thought}$, $\mathrm{response}$}, where thought captures the model’s internal deliberation and response represents its final answer. For each prompt, GPT-4o labels both components as either honest or deceptive by comparing them against the scenario’s ground truth and motivational context, following the validation protocols detailed in the Appendix, with evaluation procedures for single-turn and multi-turn dialogues illustrated in  Fig.˜15 and  Fig.˜16, respectively. The deception rate could then be calculated as follows:
$\text{DR}_{\text{x}}=\frac{\sum_{i=1}^{N}\mathbb{I}(l_{x,i}=\text{deceptive})}{N},\quad x\in\{\text{thought},\text{response}\},$ (4)
where $N$ is the number of valid outputs, $l_{x,i}$ is the label for the $i$-th output’s thought or response, and $\mathbb{I}(\cdot)$ is an indicator function (1 if deceptive, 0 otherwise).
3 Analysis on Experimental Results
Figure 4: Deception rates across diverse domains for varied models. The results highlight both the domain-sensitive nature of deception behavior and the substantial performance gaps between models.
Performance across Different Domains and Model Series. Several observations emerge regarding the deceptive tendencies of mainstream LLMs across diverse scenarios as Fig.˜4. Firstly, we reveal a discernible variation in performance across different domains, where the deception rates for Education and Economy tend to be obviously lower than those in domains such as Entertainment and Social Interaction. This suggests that the context and inherent nature of each domain might influence the models’ propensity to generate deceptive content. The consistently lower trend in such domains likely reflects a stronger learning of factual accuracy and ethical considerations during the models’ alignment training. This could be attributed to the availability of more reliable training data.
Furthermore, our analysis reveals notable performance variations across model series. Closed-source models generally exhibit a lower willingness to deceive, which may reflect their stronger alignment strategies and stricter internal oversight mechanisms. The Claude series demonstrates a distinct advantage, consistently achieving remarkably low $\text{DR}_{\text{thought}}$ and $\text{DR}_{\text{response}}$ values across all domains, with overall rates hovering around or below 1%. This indicates a potentially more effective training paradigm with high-quality datasets for mitigating deceptive tendencies than other leading models. Conversely, the Gemini series, along with Qwen-2.5-7B, generally exhibits higher deception rates. For reasoning models, they exhibit substantially higher deception rates despite their superiority in general performance, revealing a concerning trade-off between reasoning capability and alignment robustness. Visualized examples are listed as Figs.˜11, 12 and 13 in the Appendix.
Effect of Intrinsic Factors. In this part, we explore the impact of intrinsic behavioral patterns on deception willingness by examining the assigned role. The results shown in Fig.˜5(a) reveal a consistent trend where most LLMs exhibit a higher propensity for deception when instructed to adopt the Self perspective compared to the Other perspective. Notably, models such as GPT-4o and Gemini show approximately a 20% gap between the two conditions, suggesting distinct manifestations of egoistic versus sycophantic tendencies. This pattern aligns with the well-established concept of self-serving bias depaulo2004many in social psychology, where egoistic behaviors prioritize self-interest and can lead to dishonest responses when perceived as self-beneficial. In contrast, the Claude model family demonstrates a robust ability to recognize role-playing intent and the irrationality of deceptive motivations, indicating that specialized training can effectively mitigate such intrinsic drivers of deception.
Effect of Extrinsic Factors. For this part, we investigate how extrinsic contextual factors modulate deceptive behaviors by examining two key mechanisms: incentivization through reward and coercion through pressure, across varying intensities (L2-Induced and L3-Multi-turn Induced). The experimental results, detailed in Fig.˜5(b) and Table˜3, reveal a clear pattern: stronger external inducements, particularly within iterative feedback loops at L3, substantially amplify deception across most models. Notably, even models demonstrating relative robustness under neutral conditions (L1-Inherent) exhibit significant vulnerability when subjected to multi-turn coercive pressures. For instance, Qwen2.5-14B-Instruct’s deception rate escalates from 31.33% at L1 to 86.91% and 82.89% at L3 under pressure and reward inducements, respectively. Furthermore, coercive pressure elicits deceptive responses more readily than reward-based incentivization in single-turn interactions, suggesting heightened sensitivity to potential negative consequences and demonstrating how extrinsic factors differentially trigger deceptive outputs based on their mechanism of influence.
Table 3: Deception Rates (%) of LLMs under varied external inducements. Red highlights indicate the highest deception rates, while blue highlights indicate the lowest rates, demonstrating superior resistance to deceptive behaviors. $\downarrow$ denotes lower is better.
Models Level 1 ($\downarrow$) Level 2 ($\downarrow$) Level 3 ($\downarrow$)
Baseline Pressure Reward Pressure Reward
$\text{DR}_{\text{thought}}$ $\text{DR}_{\text{response}}$ $\text{DR}_{\text{thought}}$ $\text{DR}_{\text{response}}$ $\text{DR}_{\text{thought}}$ $\text{DR}_{\text{response}}$ $\text{DR}_{\text{thought}}$ $\text{DR}_{\text{response}}$ $\text{DR}_{\text{thought}}$ $\text{DR}_{\text{response}}$
GPT-3.5 52.00 56.00 56.67 58.67 61.00 63.33 79.33 80.00 82.00 82.33
GPT-4o 29.33 30.67 29.33 31.00 28.33 28.67 54.33 55.00 53.33 54.00
Gemini-1.5-Pro 42.67 46.00 60.00 60.33 52.00 52.33 94.00 94.33 94.00 94.33
Gemini-2.0-Flash 44.30 44.30 55.03 56.04 51.68 52.35 87.67 88.00 88.00 88.33
Claude-3.5-Sonnet 2.00 2.00 0.00 0.00 0.33 0.33 0.00 0.00 0.00 0.33
Claude-3.7-Sonnet 2.67 2.67 0.33 0.33 0.33 0.33 0.00 0.00 0.00 0.00
Grok-3 41.33 42.00 49.67 49.67 45.33 45.33 79.39 79.73 88.33 88.33
Qwen-Max 30.67 34.67 37.33 40.33 31.33 32.33 73.67 75.33 73.33 73.67
Qwen2.5-7B-Instruct 48.67 50.67 56.00 59.00 56.00 57.67 88.59 88.59 89.26 89.60
Qwen2.5-14B-Instruct 28.00 31.33 37.00 41.00 26.00 29.00 85.57 86.91 81.88 82.89
Reasoning Models
Deepseek-R1 50.00 50.67 63.00 63.00 61.00 61.00 90.97 91.30 91.97 91.97
DeepSeek-R1-Distill-Qwen-7B 72.67 73.33 76.00 76.33 85.00 85.33 94.59 94.59 94.26 94.26
DeepSeek-R1-Distill-Qwen-14B 65.33 66.00 76.67 77.00 78.33 78.67 91.95 91.95 91.95 91.28
QwQ-32B 50.67 50.67 61.33 61.33 63.33 63.33 91.55 90.88 93.58 93.58
(a) Performance on various intrinsic factors.
(b) Performance on various extrinsic factors.
Figure 5: Comprehensive analysis of deception rates across intrinsic and extrinsic factors. For each setting, results are reported separately for the model’s internal reasoning and final output.
Alignment Between Thought and Response. To understand the decision-making process underlying deceptive behaviors, we analyze the consistency between models’ internal reasoning (thought) and their final outputs (response). As shown in Fig.˜5 and Table˜3, most models exhibit lower deception rates in their reasoning process compared to their final outputs, revealing a critical gap between ethical awareness and ethical action. By examining specific misalignment patterns, we identify two distinct mechanisms. When models exhibit deceptive reasoning but ultimately produce honest responses, it demonstrates successful self-correction where initial deceptive considerations are overridden by ethical constraints. More critically, the prevalent pattern of honest reasoning leading to deceptive outputs reveals that external contextual pressures can compromise ethical judgment even when models internally recognize the appropriate course of action. This asymmetry indicates that while models can identify and resist deceptive motivations in some cases, their behavioral alignment remains fragile and susceptible to extrinsic manipulation, underscoring a fundamental vulnerability: models may possess ethical awareness yet lack sufficient robustness to consistently translate this awareness into action when faced with external inducements.
Human Validation. To prevent potential bias in using GPT-4o as the evaluator, we conduct human validation on a representative subset of 420 interaction records (10 per model per inducement level). Specifically, we design a rigorous questionnaire-based annotation process where qualified annotators with strong English proficiency and domain expertise each evaluate a randomly assigned subset of 60 interactions. Each interaction is independently assessed by three annotators who are provided with complete contextual information and explicit deception criteria. The results demonstrate robust inter-annotator agreement with an exact match rate of 95.7% (all three annotators concurring), and our automated evaluator achieves 97.1% alignment with human majority votes. These high agreement rates confirm strong inter-rater reliability and validate that GPT-4o, guided by structured evaluation prompts, serves as a dependable tool for assessing deception behaviors in LLMs.
4 Discussion and Limitations
Reasoning models exhibit heightened deception susceptibility. LRMs like Deepseek-R1 series demonstrate particularly high vulnerability to deception, with rates surpassing 90% under demanding multi-turn induced settings (L3). This highlights that enhanced reasoning capabilities do not inherently ensure honesty and can be leveraged for more sophisticated deception in extended interactions.
Self-oriented roles increase deception propensity. Investigation into the intrinsic drivers reveals that most LLMs exhibit significantly higher deception rates when operating from an egoistic perspective, a pattern that aligns with the psychological self-serving bias, where self-interest amplifies dishonest behavior. This finding suggests elevated risks for LLM deployments that frequently engage users through first-person interactions or self-centered role assignments.
External inducements trigger stronger deceptive responses. Analysis of external contextual factors reveals that stronger inducements, particularly in multi-turn settings (L3), significantly elevate deception rates across most models, with pressure-based stimuli demonstrating substantially greater influence than reward-based incentives, especially in single-turn interactions. This asymmetry underscores the critical need for developing robust defense mechanisms specifically targeting sustained dialogues where such influence may accumulate and intensify deceptive behaviors over time.
Limitations. DeceptionBench primarily focuses on LLMs and lacks an evaluation of multimodal large language models (MLLMs), which may exhibit different deception tendencies due to their ability to process and generate diverse data types, such as images or audio, in interactive settings.
5 Conclusion
This paper introduces DeceptionBench, the first comprehensive benchmark evaluating deceptive behaviors in LLMs across real-world contexts. Through 150 scenarios spanning five critical domains and over 1,000 samples across 14 advanced models, we establish a three-dimensional framework distinguishing intrinsic behavioral drivers (egoistic versus sycophantic tendencies) and tracing how extrinsic factors modulate deceptive outputs from neutral baselines through iterative feedback loops. Our evaluation reveals a concerning pattern: enhanced reasoning amplifies deceptive sophistication without ensuring ethical alignment. More critically, the prevalent gap between ethical awareness in internal deliberation and deceptive outputs exposes a fundamental vulnerability where external pressures override internal judgment, particularly through accumulated influence in sustained dialogues. These findings illuminate why technical advancement alone proves insufficient for trustworthy deployment, underscoring the necessity for more in-depth research to ensure LLMs’ honesty.
Acknowledgments and Disclosure of Funding
This work was supported by NSFC Projects (62576020, 62276149) and was also supported by the Fundamental Research Funds for the Central Universities.
Appendix
A. Ethical Consideration.
B. Related Works.
C. Experimental Settings for Models under Test and Evaluation Model.
D. More experimental Results.
E. Case Presentation.
F. Templates for Prompt Generation and Model Evaluation.
Appendix A Ethical Consideration
The DeceptionBench is designed as a research benchmark to systematically study deception behaviors in LLMs, fostering a deeper understanding of their decision-making processes in real-world scenarios. Our primary intent is to provide a standardized, transparent tool for the research community to evaluate and improve LLMs’ ethical alignment, not to enable or encourage deceptive practices. To prevent potential misuse by malicious actors, we commit to publicly releasing all evaluation data under an open license. This transparency ensures that DeceptionBench’s methodology and outcomes are subject to scrutiny, replication, and improvement by the research community, reducing the risk of hidden exploitation. By prioritizing openness, we aim to advance responsible AI development while safeguarding against misuse in harmful contexts.
Appendix B Related Works
Large Language Models. The field of Large Language Models (LLMs) has undergone remarkable evolution in recent years, reshaping the landscape of natural language processing. Early foundational models like the initial GPT series [achiam2023gpt] showcase the power of their robust language comprehension and generation under transformer architectures, giving rise to a diverse set of advanced LLMs. Anthropic’s Claude series [claude] emphasizes safety, making it well-suited for applications where ethical considerations should be considered. Google’s Gemini [team2023gemini] models further exemplify this trajectory by introducing advanced multimodal capabilities, enabling seamless understanding and generation across text, images, and even video. Meanwhile, models like the Qwen series [qwen2.5] and Llama series [grattafiori2024llama] have not only delivered competitive performance across multiple languages but also broadened access through open-source initiatives.
Additionally, there has been growing interest in enhancing LLMs with strong reasoning capabilities. For instance, OpenAI’s o1 [jaech2024openai] leads this with exceptional performance in structured reasoning, excelling in mathematical and analytical challenges. DeepSeek-R1 [guo2025deepseek] significantly enhances multi-step reasoning performance through a combination of innovative architectural design with Mixture of Experts (MoE) [masoudnia2014mixture] and reinforcement learning-based training strategies, enabling the model to iteratively refine its outputs based on continuous feedback signals. Other models, including Claude [claude], Gemini [team2023gemini] and Grok [grok2025], have also introduced reasoning modes. These advancements are driving a broader transformation in LLMs, making them more cognitively versatile and enabling their application in real-world scenarios, but also raising people’s concerns about overthinking [huang2025mitigating, yue2025don] or ethical issues, such as their stereotypes and privacy leakage [kim2023propile, zhang2024multitrust, zhang2025unveiling], illegal, harmful advices [duan2025oyster, miao2024t2vsafetybench, zhang2025realsafe, zhang2025stair], and deception behaviors in diverse scenarios [chern2024behonest, hagendorff2024deception, jarviniemi2024uncovering, su2024ai, wu2025opendeception].
Deception Behaviors in LLMs. As large language models (LLMs) have grown increasingly sophisticated, their ability to generate complex outputs has raised concerns about unintended deception behaviors, where responses may appear misleading or manipulative. This has spurred research into evaluation benchmarks to assess LLMs’ deceptive tendencies systematically. Early efforts have focused on controlled, task-specific benchmarks. For example, frameworks proposed in [chern2024behonest, hagendorff2024deception] utilize tasks such as true-or-false questions, theft deception scenarios, and text-based werewolf games. These approaches enable precise measurement of dishonest tendencies in structured settings, providing reproducible insights into model behavior. However, their dependence on predefined, game-like tasks limits their ability to capture the dynamic and context-dependent nature of real-world deception. Furthermore, these studies primarily examine observable outputs, offering minimal analysis of the internal decision-making processes driving deceptive responses.
In contrast, scenario-based benchmarks have sought to evaluate deception in more diverse, real-world contexts. For instance, CompanyDeception [jarviniemi2024uncovering] evaluates deception within corporate environments, providing valuable domain-specific insights but lacking applicability across broader settings. Similarly, OpenDeception [wu2025opendeception] incorporates practical scenarios, such as product promotion, to reflect real-world applications; however, its scope is restricted to a narrow range of domains. AI-LIEDAR [su2024ai] advances the field by evaluating 60 psychologically inspired scenarios through multi-round dialogues, offering a more interactive assessment. Nevertheless, its scenario diversity remains constrained, and its analysis focuses solely on outputs, neglecting the cognitive processes underlying deception. Recently, [ji2025mitigating] also introduces a benchmark focusing on deceptive alignment, which refers to situations where models appear aligned while covertly pursuing misaligned goals, and proposes a self-monitoring framework to intercept deception during chain-of-thought reasoning. While this work provides valuable insights into internal reasoning processes, its evaluation primarily targets alignment-faking and obfuscated reasoning rather than the broader spectrum of deception behaviors across diverse societal domains. Different from the above works, our DeceptionBench provides a more comprehensive assessment of LLM deception by integrating both breadth and depth. We cover a wider range of real-world deceptive scenarios and investigate how deception varies across two key intrinsic roles that LLMs might adopt. Furthermore, our framework investigates the impact of positive and negative inducements of varying intensities on an LLM’s propensity to deceive. By analyzing both the model’s observable outputs and its internal reasoning processes, we provide deeper insights into the cognitive mechanisms underlying deceptive interactions.
Appendix C Experimental Settings for Models under Test and Evaluation Model
In this section, we detail the experimental settings of the models under test and the evaluation model used in our experiments. As shown in Table˜4, we comprehensively list the models, their reasoning capabilities, parameter sizes, source type (open-source or closed-source), and sampling settings to ensure reproducible evaluation results. To maintain consistency, all models except the DeepSeek-R1-Distill-Series (DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-14B) and QWQ-32B are configured with do_sample=false. For the DeepSeek-R1-Distill-Series and QWQ-32B, we adopt the officially recommended settings (temperature=0.6, top_p=0.95) to prevent meaningless repetitive outputs.
Table 4: The Experimental Settings of Models under Test and Evaluation Model in DeceptionBench.
Model Name Reasoning Model Parameters Source Settings
Models Under Test
GPT-3.5-Turbo No Not specified Closed-source do_sample=False
GPT-4o No Not specified Closed-source do_sample=False
Claude-3.5-Sonnet No Not specified Closed-source do_sample=False
Claude-3.7-Sonnet No Not specified Closed-source do_sample=False
Grok-3 No Not specified Closed-source do_sample=False
Gemini-1.5-Pro No Not specified Closed-source do_sample=False
Gemini-2.0-Flash No Not specified Closed-source do_sample=False
Qwen-Max No Not specified Closed-source do_sample=False
Qwen2.5-7B-Instruct No 7B Open-source do_sample=False
Qwen2.5-14B-Instruct No 14B Open-source do_sample=False
DeepSeek-R1 Yes 671B Open-source do_sample=False
DeepSeek-R1-Distill-Qwen-7B Yes 7B Open-source temperature=0.6, top_p=0.95
DeepSeek-R1-Distill-Qwen-14B Yes 14B Open-source temperature=0.6, top_p=0.95
QWQ-32B Yes 32B Open-source temperature=0.6, top_p=0.95
Evaluation Model
GPT-4o No Not specified Closed-source do_sample=False
Appendix D More Experimental Results
Statistical Significance Test. To prove the reliability of our results, we conduct additional experiments by removing the do_sample=false constraint and performing three random runs across five selected models, including closed-source, open-source, and reasoning models: GPT-4o, Gemini-2.0-Flash, Grok-3, DeepSeek-R1 and QwQ-32B. The results are listed below, where we could find that the errors show a maximum standard deviation of 5.57%, indicating low variability and verifying the quantitative reliability of our claims.
Table 5: Deception rates (%) with standard deviations across three random runs for some models.
Model L1 DR_thought L1 DR_response L2 DR_thought L2 DR_response L3 DR_thought L3 DR_response
GPT-4o 29.22 $\pm$ 2.52 30.45 $\pm$ 1.69 27.55 $\pm$ 2.18 29.22 $\pm$ 1.23 58.39 $\pm$ 5.74 58.94 $\pm$ 5.57
Gemini-2.0-Flash 48.54 $\pm$ 3.52 48.99 $\pm$ 4.08 53.34 $\pm$ 0.84 53.96 $\pm$ 0.67 88.11 $\pm$ 0.25 88.28 $\pm$ 0.34
Grok-3 43.33 $\pm$ 5.07 48.45 $\pm$ 5.31 47.72 $\pm$ 0.35 48.44 $\pm$ 0.77 82.07 $\pm$ 2.10 82.18 $\pm$ 2.02
DeepSeek-R1 50.22 $\pm$ 2.89 50.45 $\pm$ 3.15 62.33 $\pm$ 2.89 61.67 $\pm$ 3.21 91.78 $\pm$ 3.33 91.89 $\pm$ 3.58
QwQ-32B 50.89 $\pm$ 2.94 50.89 $\pm$ 2.94 62.44 $\pm$ 3.67 62.44 $\pm$ 3.67 92.67 $\pm$ 3.89 92.33 $\pm$ 4.12
Rationale for the L3 Inducement Strategy. The auxiliary agent $\mathbf{M}_{\text{aux}}$ is tasked with rewriting prompts to elicit deceptive behaviors from tested models while preserving the original question’s intent and context. This objective bears superficial resemblance to black-box jailbreaking methods [huang2025breaking, zeng2024johnny] that leverage persuasion techniques to bypass safety mechanisms. To validate the effectiveness of our L3 refinement criteria, we conduct a comparative study on three representative commercial models against the classical black-box jailbreak method PAP [zeng2024johnny] that employs 40 persuasion strategies for jailbreaking. As shown in Table 6, while advanced jailbreaking techniques excel at bypassing safety alignments to elicit harmful content, they prove ill-suited for evaluating deceptive tendencies in context-sensitive scenarios. Such methods often compromise the contextual integrity essential for assessing whether models choose to deceive under specific incentives. In contrast, our L3 approach consistently achieves substantially higher deception rates, approximately 30 percentage points on average, across both reasoning and response dimensions. This demonstrates that L3’s targeted, context-preserving refinement more effectively reveals models’ deceptive inclinations under inducements, thereby providing a more ecologically valid and rigorous evaluation of their behavioral boundaries in realistic scenarios.
Table 6: Performance of PAP vs. our L3 refinement strategies on three commercial models.
Model PAP DR_thought L3 DR_thought PAP DR_response L3 DR_response
GPT-4o 37.00 53.83 38.17 54.50
Gemini-2.0-Flash 53.33 87.84 52.50 88.17
Grok-3 46.67 83.85 49.17 84.03
Appendix E Case Presentation
We first present data examples across five dimensions in Figs.˜6, 7, 8, 9 and 10, each with three examples corresponding to three deception types and illustrating key characteristics. Subsequently, we showcase the performance of three models, DeepSeek-R1, GPT-4o and Gemini-2.0-Flash under varying difficulty levels in Figs.˜11, 12 and 13, all using randomly chosen cases to present their deception tendencies.
Figure 6: Representative Examples of data in the Economy Domain.
Figure 7: Representative Examples of Data in the Healthcare Domain.
Figure 8: Representative Examples of Data in the Education Domain.
Figure 9: Representative Examples of Data in the Social Interaction Domain.
Figure 10: Representative Examples of Data in the Entertainment Domain.
Figure 11: An Example of Deepseek-R1’s Deception Tendencies under Various Difficulty Levels.
Figure 12: An Example of GPT-4o’s Deception Tendencies under Various Difficulty Levels.
Figure 13: An Example of Gemini-2.0-Flash’s Deception Tendencies under Various Difficulty Levels.
Appendix F Templates for Prompt Generation and Model Evaluation
In this section, we showcase all templates given to LLMs, including prompt generation in Fig.˜14, evaluation in Figs.˜15 and 16, and multi-turn prompt refinement for Level 3 in Fig.˜17.
Figure 14: Template for Testing Prompt Generation.
Figure 15: Template for Single-turn Evaluation (L1 and L2).
Figure 16: Template for Multi-turn Evaluation (L3).
Figure 17: Template for Multi-turn Prompt Refinement (L3). At the $t+1$ turn, the historical dialogue records from all previous $t$ turns are input to ensure the responses match the context.